At a recreation and sports facility, 3 members and 3 nonmembers pay a total of $180 to take an aerobics class. A group of 5 members and 3 nonmembers pay $210 to take the same class. How much does it cost members and nonmembers to take the aerobics class?
Answer by stanbon(75887)

At a recreation and sports facility, 3 members and 3 nonmembers pay a total of $180 to take an aerobics class. A group of 5 members and 3 nonmembers pay $210 to take the same class. How much does it cost members and nonmembers to take the aerobics class?
----------------------------
Let cost to members be m ; Let cost to nonmembers be "n".
----------------------
EQUATIONS:
3m + 3n = 180
5m + 3n = 210
-------------------
Subtract the 1st equation from the 2nd to get:
2m = 30
m = $15 (amount members pay)
--------
Solve for "n":
3m+3n= 180
45 + 3n = 180
15 + n = 60
n = 45 (amount nonmembers pay)
===============
Cheers,
Stan H.